OpenStudy (kodak.kiid):

write 4 7/10 as a decimal number.

OpenStudy (qwertty123):

To get a decimal approximation of the fraction, simply divide the fraction's numerator by its denominator.

OpenStudy (qwertty123):

7 / 10 = ?

OpenStudy (thedolorosablu3capricousg1rl):

knowing this, I'm sure it would be 4.7. Either way Qwertty is right with the process of it

OpenStudy (kodak.kiid):

i found it out thank you its 0.7

OpenStudy (qwertty123):

Now add 0.7 to the whole number portion of the mixed number. 0.7 + 4 =

OpenStudy (qwertty123):

@Kodak.Kiid Hello... 0.7+4=?

OpenStudy (kodak.kiid):

4.7

OpenStudy (kodak.kiid):

@Qwertty123

OpenStudy (qwertty123):

Correct! Nice Job ^.^

OpenStudy (kodak.kiid):

thank you @Qwertty123

OpenStudy (qwertty123):

You Welcome! ^,^